# Approximate Bayesian Computation sequential Monte Carlo via random forests

### v1

@@ 8                  If $\pi\left(\theta_{t}^{(i)}\right)=0$, then return to Step 6
We have developed an R package abc-smc-rf based on Algorithms 5 and 6. It offers several practical advantages when compared to direct applications of abcrf and drf. First, constructing iterative random forests of size $B$ from reference tables of sizes $N_{1},\dots,N_{T}$ requires $\mathcal{O}\left(B\cdot\sum_{t=1}^{T}(N_{t}\cdot\log N_{t})\right)$ splits, assuming balanced trees. This requires less computational resources than growing a random forest based on a reference table of size $N=\sum_{i=1}^{T}N_{i}$, which contains $\mathcal{O}\left(B\cdot N\log N)\right)$ splits.
Therefore, with the same total number of model simulations, ABC-SMC-(D)RF requires less memory and computational runtime. Second, similar to the behavior of ABC-SMC compared to ABC-REJ, ABC-SMC-(D)RF is likely to converge to the posterior distribution faster than ABC-RF or ABC-DRF, because the parameter distributions are constantly updated to focus on regions in the parameter space $\Theta$ that best explain $s^{obs}$. We will test this hypothesis with some examples in the next section. Finally, unlike ABC-RF and ABC-DRF which require the full reference table before forest construction and parameter prediction, the approximated posterior distribution from each iteration of ABC-SMC-(D)RF can be compared against the previous iteration to assess whether it has converged, potentially further lowering computational expense.
@@ 5 Results
In this section, we compare the performance of ABC-SMC-(D)RF (Algorithm 5 or 6) across different inference problems, against ABC-REJ (Algorithm 7), MCMC (Algorithm 8 or 9), ABC-SMC (Algorithm 10), ABC-RF (Algorithms 1, 2) and ABC-DRF (Algorithms 3, 4). Unless specified otherwise, for ABC-RF, ABC-DRF and ABC-SMC-(D)RF, we use the default parameters in abcrf and drf for tree count $B$, candidate statistic count $n_{try}$, leaf size threshold $N_{min}$, root node subsample size $n_{sample}$, etc.
@@ 5.1 Deterministic Lotka-Volterra model
where $a$ is the prey’s birth rate, $d$ is the predator’s death rate, and $c$ and $b$ are the prey’s death rate and predator’s birth rate due to predation, respectively. We fix $c=d=1$ and seek to infer $\theta=(a,b)$ from prior distribution $a,b\sim\text{Uniform}(-10,10)$. Similar to Toni et al. (2008), we solve the ODE system with initial condition $(x(0),y(0))=(1,0.5)$ with $\theta=(1,1)$, and sample $(x,y)$ at eight time points between $t=0$ and $t=15$, then add Normal$(0,0.5^{2})$ noise to each data point to form the observed statistics $s^{obs}=\left\{x_{1},y_{1},\dots,x_{8},y_{8}\right\}$.
We first implement ABC-SMC-(D)RF for multiple parameters. Algorithm 6 is applied with $T=4$ iterations, each with $N_{t}=5,000$ simulations in the reference table, and perturbation kernels $K_{t}\left(\theta\middle|\theta^{*}\right)=\theta^{*}+\text{Uniform}(-0.1,0.1)$. Figures 3a, b show the distributions of $a$ and $b$ after each iteration $t$ in abc-smc-rf. The distributions become more concentrated as $t$ progresses, but stay centered around the true values of $(a,b)$ (Table 1).
To evaluate abc-smc-rf’s performance, we compare the final posterior distributions with ABC-DRF and ABC-SMC. Algorithm drf is performed on a reference table of size $N=20,000$, to match the total number of simulations in abc-smc-rf. We use the R package EasyABC (Jabot et al., 2023) to implement ABC-SMC with $T=5$ iterations. EasyABC uses the distance $d\left(s,s^{obs}\right)=\sum_{i=1}^{8}\left[(x_{i}-x_{i}^{obs})^{2}+(y_{i}-y_{i}^{obs})^{2}\right]$ to compare simulated statistics $s$ against observation $s^{obs}$, and adaptive perturbation kernels $K_{t}\left(\theta\middle|\theta^{\prime}\right)=\text{Normal}\left(\theta^{\prime},2\cdot\text{var}\left(\theta_{t-1}\right)\right)$ proposed in Beaumont et al. (2009). Referencing Toni et al. (2008), we impose the series of tolerance thresholds $\epsilon_{1}=30.0,\epsilon_{2}=16.0,\epsilon_{3}=6.0,\epsilon_{4}=5.0,\epsilon_{5}=4.3$. To produce 1,000 accepted particles, EasyABC requires $N=56,850$ simulations.
Figures 3c, d present the posterior distributions approximated by the ABC implementations. The marginal distributions from ABC-SMC-DRF are as centered and concentrated around the true values for $(a,b)$ as ABC-SMC (Table 1). In contrast, even though the distributions from DRF are centered around the ground truth, the variances are consistently higher, indicating more uncertainty.
@@
5.2 Linear birth-death process
We next examine ABC-SMC-DRF’s performance for the linear birth-death process example in Tavaré (2018). The model starts at time 0 with $Z(0)$ individuals, each of which divides at rate $\lambda>0$ and dies at rate $\mu>0$. The number of individuals at time $t>0$ is denoted $Z(t)$, with expected value $\mathbb{E}Z(t)=Z(0)\mathrm{e}^{(\lambda-\mu)t}$.
The probability $\mathbb{P}\left(Z(t_{i+1})=z_{i+1}\middle|Z(t_{i})=z_{i},\lambda,\mu\right)$ that there are $z_{i+1}$ individuals at time $t_{i+1}$ given that there are $z_{i}$ individuals at time $t_{i}$ follows from Keiding (1975) as
$\sum_{l=0}^{\min(z_{i},z_{i+1})}\binom{z_{i}}{l}(1-\alpha(t))^{l}\alpha(t)^{z_{i}-l}\cdot\binom{z_{i+1}-1}{z_{i+1}-l}(1-\beta(t))^{l}\beta(t)^{z_{i+1}-l}$
where $\alpha(t)$ and $\beta(t)$ are defined by
[…]
(Kendall, 1948). The Markov property implies that the likelihood for $\lambda$ and $\mu$, given an observation $y^{obs}$ where $Z(t_{1})=z_{1},\dots,Z(t_{n})=z_{n}$, is:
$f\left(\lambda,\mu\middle|y^{obs}\right)=\prod_{i=0}^{n-1}\mathbb{P}\left(Z(t_{i+1})=z_{i+1}\middle|Z(t_{i})=z_{i},\lambda,\mu\right)$ (9)
Our inference problem consists of approximating the posterior distribution of $\theta=(\lambda,\mu)$ from a prior distribution $\lambda,\mu\sim$ Uniform$(0,20)$, for observed data $s^{obs}=y^{obs}=(z_{i},t_{i})_{i=1,\dots,25}$ from initial condition $Z(0)=10$ in (Tavaré, 2018, Table 3). The true posterior distribution can be computed from (9). Similarly, we build a MCMC chain using the Metropolis-Hastings algorithm 8 with likelihood function (9) and uniform proposal kernel $q\left(\lambda,\mu\middle|\lambda^{\prime},\mu^{\prime}\right)=(\lambda^{\prime}+\text{U}(-1,1),\mu^{\prime}+\text{U}(-1,1))$ to approximate the posterior distribution. After 2500 burn-in steps, we accepted 1950 simulations from a total of $N=100,000$ steps (average acceptance rate = 0.379) spaced every 50 steps. Finally, we implement drf with a reference table of size $N=20,000$, abcrf separately for $\lambda$ and $\mu$ likewise with $N=20,000$, and abc-smc-rf using the multivariate option with $T=4$ iterations, each with $N_{t}=5,000$ simulations. All RF methods approximate the posterior distributions from $B=2,500$ trees. ABC-SMC-DRF implements perturbation kernels $K_{t}(\theta|\theta^{\prime})=\theta^{\prime}+\text{U}(-2,2)\times\text{U}(-2,2)$.
Unsurprisingly, compared to the true density, MCMC is the most accurate in its inferred joint and marginal posterior distributions (Figure 4), as it utilizes the likelihood function in accepting new particles in the chain. The joint distributions from ABC-DRF and ABC-SMC-DRF are concentrated in the same area as MCMC, and exhibit a similar correlation pattern between $\lambda$ and $\mu$ (Figure 4a). However, quantile-quantile plots comparing the marginal distributions of $\lambda$ and $\mu$ between the exact posterior and the estimates reveal that ABC-DRF’s result has a significant bias toward larger values, compared to MCMC (Figure 4b, c). The marginal distributions from ABC-RF have reduced tails and are closer to the true density. However, the results from ABC-SMC-DRF are the most accurate among the RF implementations, and are almost as good as those from MCMC (Figure 4b, c).
Figure 4: Parameter inference for the linear birth-death branching process. a: Joint posterior distributions for birth rate $\lambda$ and death rate $\mu$, computed with the likelihood (Eq. 9, density heatmap) and estimated with MCMC, ABC-DRF with CART splitting rule, and ABC-SMC-DRF. b, c: Marginal quantile-quantile plots for inferred $\lambda$ (b) and $\mu$ (c) from MCMC, ABC-RF, ABC-DRF, and ABC-SMC-DRF (y axis), against the exact likelihood-based distributions (x axis).
@@ 5.3 Stochastic biochemical reaction systems
Finally, we examine ABC-SMC-DRF’s application in inferring reaction rates for biochemical systems. We first examine the Michaelis-Menten model, which describes the kinetics between an enzyme $E$ that binds to a substrate $S$ to form a complex $ES$, leading to product $P$:
@@
respectively, where the Avagadro’s constant nA $=6.023\cdot 10^{23}$ approximates the number of molecules in a mole, vol $=10^{-15}$ is the volume of the system, and $\theta=(c_{1},c_{2},c_{3})$ parameterizes the Michaelis-Menten model for different biochemical systems (Wilkinson, 2018, Example 7.3).
The initial state is defined by $E(0)=2\cdot 10^{-7}\cdot\text{nA}\cdot\text{vol}$, $S(0)=5\cdot 10^{-7}\cdot\text{nA}\cdot\text{vol}$, and $ES(0)=P(0)=0$ (Wilkinson, 2018). We seek to infer $\theta$ from $s^{obs}=\left\{X(t),t=1,\dots,10\right\}$, simulated with true parameters $c_{1}=0.1,c_{2}=6,c_{3}=-4$. We compare the results from drf with $N=$ 20,000 simulations, and abc-smc-rf with $T=$ 5 iterations, each with reference tables of size $N_{t}=$ 4,000, from uniform prior distributions $\pi\left(c_{1}\right)=$ U$(0,1)$, $\pi\left(c_{2}\right)=$ U$(5,7)$, $\pi\left(c_{3}\right)=$ U$(-5,-3)$. The perturbation kernels in abc-smc-rf in each iteration are $K_{t}\left(c_{1}\middle|c_{1}^{\prime}\right)=c_{1}^{\prime}+\text{U}(-0.05,0.05)$, $K_{t}\left(c_{2}\middle|c_{2}^{\prime}\right)=c_{2}^{\prime}+\text{U}(-0.1,0.1)$, and $K_{t}\left(c_{3}\middle|c_{3}^{\prime}\right)=c_{3}^{\prime}+\text{U}(-0.1,0.1)$. All simulations are generated by applying Gillespie’s algorithm (Gillespie, 1977).
Comparing the results for $c_{1}$ and $c_{2}$ shows that ABC-SMC-DRF’s posterior distributions ($c_{1}$: mean = 0.110, 95% CI = $[0.083,0.141]$; $c_{2}$: mean = 5.959, 95% CI = $[5.905,6.016]$) are centered closely around the true values of $c_{1}=0.1,c_{2}=6$, similarly to ABC-DRF, but with narrower confidence intervals (ABC-DRF $c_{1}$: mean = 0.102, 95% CI = $[0.023,0.176]$; $c_{2}$: mean = 5.943, 95% CI = $[5.759,6.093]$) (Figure 5a, b). In contrast, both methods are uncertain about the true distribution for $c_{3}$ (ABC-SMC-DRF: mean = -3.994, 95% CI = $[-4.887,-3.259]$; ABC-DRF: mean = -3.950, 95% CI = $[-4.915,-3.032]$) (Figure 5c).
@@
Indeed, despite the uncertainty in $c_{3}$, simulations performed with $\theta$-values sampled from the posterior distributions of ABC-DRF and ABC-SMC-DRF are centered around the observed values $s^{obs}$ across different molecules and time points (Figure 5e, f, g, h). Moreover, consistent with ABC-SMC-DRF’s higher certainty in the inference for $c_{1}$ and $c_{2}$, molecule counts simulated from its posterior distribution have significantly reduced range compared to ABC-DRF.

### v2

@@ 19  Normalize $\left\{w_{t}^{*(i)}\right\}$
The second version, ABC-SMC-RF (Algorithm 7), is built around ABC-RF, and designed for univariate problems or finding marginal distributions for multivariate models. Each parameter in an iteration is sampled independently from the previous iteration and then perturbed. The statistics are then computed for the whole parameter set. The reference table consists of these statistics and values corresponding to a given parameter, from which ABC-RF predicts the parameter’s marginal distribution based on $s^{\rm obs}$. The ABC-RF weights are then re-calibrated and normalized.
1 for $t=1,\ldots,T$ do
2 for $i=1,\dots,N_{t}$ do
3    if $t=1$ then
4         Generate $\theta_{1}^{(i)}\sim\pi(\theta)$
5    else
6       for $j=1,\dots,|\Theta|$ do
7            Sample $\theta^{*}_{j}$ from $\left\{\left(\theta^{(k)}_{j,t-1},w^{*(k)}_{j,t-1}\right)\right\}_{k=1,\dots,N_{t-1}}$
8           Generate $\theta_{j,t}^{(i)}\sim K_{j,t}(\theta|\theta_{j}^{*})$
9           If $\pi_{j}\left(\theta_{j,t}^{(i)}\right)=0$, then return to Step 7
10
11       $\theta_{t}^{(i)}\leftarrow\left(\theta_{1,t}^{(i)},\dots,\theta_{|\Theta|,t}^{(i)}\right)$
12
14
15 for $j=1,\dots,|\Theta|$ do
17     Perform ABC-RF (Algorithms 1 and 2) with reference table $R_{j}$ to compute weights $w_{j,t}^{(1)},\dots,w_{j,t}^{(N_{t})}$ for observed statistics $s^{\rm obs}$
18    if $t=1$ then
19       $w_{j,t}^{*(i)}=w_{j,t}^{(i)}$
20    else
21       $w_{j,t}^{*(i)}=w_{j,t}^{(i)}\frac{\pi_{j}\left(\theta_{j,t}^{(i)}\right)}{\sum_{k=1}^{N_{t-1}}w_{j,t-1}^{*(k)}\cdot K_{j,t}\left(\theta_{j,t}^{(i)}\middle|\theta_{j,t-1}^{(k)}\right)}$
22
23     Normalize $\left\{w_{j,t}^{*(i)}\right\}$
24
Algorithm 7 ABC-SMC-RF for single parameters
The kernel values $H_{h}\left(\left\|s-s^{\rm obs}\right\|\right)$ in ABC-SMC (Algorithm 5) increase as $s\rightarrow s^{\rm obs}$. In the ABC-SMC-(D)RF algorithms, these kernels are replaced by the weights from ABC-DRF in Algorithm 6 and ABC-RF in Algorithm 7. These weights behave similarly to the ABC-SMC kernels: particles with statistics $s$ closer to $s^{\rm obs}$ fall into the same leaves as the data more often across the forest, therefore the corresponding parameters $\theta$ are more heavily weighted. Furthermore, in ABC-SMC, smaller scale parameters $h_{t}$ reduce the kernel diffusion, that is, only particles with $\left\|s-s^{\rm obs}\right\|$ increasingly close to 0 have high weights. Decreasing the minimum node sizes $N_{min}$ in ABC-(D)RF has the same effect, as it reduces the accepted particles across the forest to only those with statistics very close to those of the data. Finally, ABC-SMC-(D)RF does not reject particles before (D)RF is performed in each iteration except for those outside the prior distribution, and hence is analogous to ABC-SMC with rejection thresholds $c_{2}=\dots=c_{T}=0$.
We have developed an R package abcsmcrf based on Algorithms 6 and 7. It offers several practical advantages when compared to direct applications of abcrf and drf. First, constructing iterative random forests of size $B$ from reference tables of sizes $N_{1},\dots,N_{T}$ requires $\mathcal{O}\left(B\cdot\sum_{t=1}^{T}(N_{t}\cdot\log N_{t})\right)$ splits, assuming balanced trees. This requires less computational resources than growing a random forest based on a reference table of size $N=\sum_{i=1}^{T}N_{i}$, which contains $\mathcal{O}\left(B\cdot N\cdot\log N)\right)$ splits. Therefore, with the same total number of model simulations, ABC-SMC-(D)RF requires less memory and computational runtime. Second, similar to the behavior of ABC-SMC compared to ABC-REJ, ABC-SMC-(D)RF is likely to converge to the posterior distribution faster than ABC-RF or ABC-DRF, because the parameter distributions are constantly updated to focus on regions in the parameter space $\Theta$ that best explain $s^{\rm obs}$. We will test this with some examples in the next section. Finally, unlike ABC-RF and ABC-DRF which require the full reference table before forest construction and parameter prediction, the approximated posterior distribution from each iteration of ABC-SMC-(D)RF can be compared against the previous iteration to assess whether it has converged, potentially further lowering computational expense.
@@ 5 Results
In this section, we compare the performance of ABC-SMC-(D)RF (Algorithm 6 or 7) across different inference problems, against ABC-SMC (Algorithm 5), ABC-RF (Algorithms 1, 2) and ABC-DRF (Algorithms 3, 4). Unless specified otherwise, for ABC-RF, ABC-DRF and ABC-SMC-(D)RF, we use the default parameters in abcrf and drf for tree count $B$, candidate statistic count $n_{\rm try}$, leaf size threshold $N_{\rm min}$, root node subsample size $n_{\rm sample}$, etc.
@@ 5.1 Deterministic Lotka-Volterra model
[…]
We first implement ABC-SMC-(D)RF for multiple parameters. Algorithm 6 is applied with $T=4$ iterations, each with $N_{t}=5,000$ simulations in the reference table, and perturbation kernels $K_{t}\left(\theta\middle|\theta^{*}\right)=\theta^{*}+\text{Uniform}(-0.1,0.1)$. Figures 2a, b show the distributions of $a$ and $b$ after each iteration $t$ in abcsmcrf. The distributions become more concentrated as $t$ progresses, but stay centered around the true values of $(a,b)$ (Table 1).
To evaluate abcsmcrf’s performance, we compare the final posterior distributions with ABC-DRF and ABC-SMC. Algorithm drf is performed on a reference table of size $N=20,000$, to match the total number of simulations in abcsmcrf. We use the R package EasyABC (Jabot et al., 2023) to implement ABC-SMC with $T=5$ iterations. EasyABC uses the distance $d\left(s,s^{\rm obs}\right)=\sum_{i=1}^{8}\left[(x_{i}-x_{i}^{\rm obs})^{2}+(y_{i}-y_{i}^{\rm obs})^{2}\right]$ to compare simulated statistics $s$ against observation $s^{\rm obs}$, and adaptive perturbation kernels $K_{t}\left(\theta\middle|\theta^{\prime}\right)=\text{Normal}\left(\theta^{\prime},2\cdot\text{var}\left(\theta_{t-1}\right)\right)$ as proposed in Beaumont et al. (2009). Referencing Toni et al. (2008), we impose the series of tolerance thresholds $\epsilon_{1}=30.0,\epsilon_{2}=16.0,\epsilon_{3}=6.0,\epsilon_{4}=5.0,\epsilon_{5}=4.3$. To produce 1,000 accepted particles, EasyABC requires $N=56,850$ simulations.
Figures 2c, d present the posterior distributions approximated by the ABC implementations. The marginal distributions from ABC-SMC-DRF are as centered and concentrated around the true values for $(a,b)$ as ABC-SMC (Table 1). In contrast, even though the distributions from DRF are centered around the ground truth, the variances are consistently higher, indicating higher uncertainty.
@@ 5.2 Stochastic biochemical reaction systems
We examine ABC-SMC-DRF’s application in inferring reaction rates for biochemical systems. We first examine the Michaelis-Menten model, which describes the kinetics between an enzyme $E$ that binds to a substrate $S$ to form a complex $ES$, leading to product $P$:
@@
respectively, where the Avagadro’s constant nA $6.023\cdot 10^{23}$ approximates the number of molecules in a mole, vol $10^{-15}$ is the volume of the system, and $\theta=(c_{1},c_{2},c_{3})$ parameterizes the Michaelis-Menten model for different biochemical systems (Wilkinson, 2018, Example 7.3).
The initial state is defined by $E(0)=2\cdot 10^{-7}\cdot\text{nA}\cdot\text{vol}$, $S(0)=5\cdot 10^{-7}\cdot\text{nA}\cdot\text{vol}$, and $ES(0)=P(0)=0$ (Wilkinson, 2018). We seek to infer $\theta$ from $s^{\rm obs}=\left\{X(t),t=1,\dots,10\right\}$, simulated with true parameters $c_{1}=0.1,c_{2}=6,c_{3}=-4$. We compare the results from drf with $N=$ 20,000 simulations, and abcsmcrf with $T=$ 5 iterations, each with reference tables of size $N_{t}=$ 4,000, from uniform prior distributions $\pi\left(c_{1}\right)=$ U(0,1)(0,1), $\pi\left(c_{2}\right)=$ U(5,7)(5,7), $\pi\left(c_{3}\right)=$ U(5,3)(-5,-3). The perturbation kernels in abcsmcrf in each iteration are $K_{t}\left(c_{1}\middle|c_{1}^{\prime}\right)=c_{1}^{\prime}+\text{U}(-0.05,0.05)$, $K_{t}\left(c_{2}\middle|c_{2}^{\prime}\right)=c_{2}^{\prime}+\text{U}(-0.1,0.1)$, and $K_{t}\left(c_{3}\middle|c_{3}^{\prime}\right)=c_{3}^{\prime}+\text{U}(-0.1,0.1)$. All simulations are generated by applying Gillespie’s algorithm (Gillespie, 1977).
Comparing the results for $c_{1}$ and $c_{2}$ shows that ABC-SMC-DRF’s posterior distributions ($c_{1}$: mean = 0.110, 95% CI = $[0.083,0.141]$; $c_{2}$: mean = 5.959, 95% CI = $[5.905,6.016]$) are centered closely around the true values of $c_{1}=0.1,c_{2}=6$, similarly to ABC-DRF, but with narrower confidence intervals (ABC-DRF $c_{1}$: mean = 0.102, 95% CI = $[0.023,0.176]$; $c_{2}$: mean = 5.943, 95% CI = $[5.759,6.093]$) (Figure 3a, b). In contrast, both methods are uncertain about the true distribution for $c_{3}$ (ABC-SMC-DRF: mean = -3.994, 95% CI = [4.887,3.259][-4.887,-3.259]; ABC-DRF: mean = -3.950, 95% CI = [4.915,3.032][-4.915,-3.032]) (Figure 3c).
@@
Indeed, despite the uncertainty in $c_{3}$, simulations performed with $\theta$ values sampled from the posterior distributions of ABC-DRF and ABC-SMC-DRF are centered around the observed values $s^{obs}$ across different molecules and time points (Figure 3e, f, g, h). Moreover, consistent with ABC-SMC-DRF’s higher certainty in the inference for $c_{1}$ and $c_{2}$, molecule counts simulated from its posterior distribution have significantly reduced range compared to ABC-DRF.